OpenStudy (anonymous):

What is the function rule for the graph shown? (see reply bar for graph and multiple choice)

OpenStudy (anonymous):

A. y = 9x B.y = x + 9 C. y = x – 9 D. \[y=\frac{ 1 }{ 9 } x\]

OpenStudy (anonymous):

That is a line. Its equation in the format of y = mx + n as you se 1) when x = 9 y= 0 so 0=m*9+n --> n = -9m 2) when x=11 y = 2 so 2 = m*11 + n put n = -9m --> 2 = 11m -9m --> 2 = 2m --> m = 1 Put m in equation 1 or 2. I'll put in 2. 2) --> 2 = 1*11 + n --> n = -9 so y = mx + n = 1x -9 = x-9 ---> y= x-9

OpenStudy (anonymous):

thank you!